# Bi-CoG: Bi-Consistency-Guided Self-Training for Vision-Language Models

### v1

@@ Abstract
Exploiting unlabeled data through semi-supervised learning (SSL) or leveraging pre-trained models via fine-tuning are two prevailing paradigms for addressing label-scarce scenarios. Recently, growing attention has been given to combining fine-tuning of pre-trained vision-language models (VLMs) with SSL, forming the emerging paradigm of semi-supervised fine-tuning. However, existing methods often suffer from model bias and hyperparameter sensitivity, due to reliance on prediction consistency or pre-defined confidence thresholds. To address these limitations, we propose a simple yet effective plug-and-play methodology named Bi-Consistency-Guided Self-Training (Bi-CoG), which assigns high-quality and low-bias pseudo-labels, by simultaneously exploiting inter-model and intra-model consistency, along with an error-aware dynamic pseudo-label assignment strategy. Both theoretical analysis and extensive experiments over 14 datasets demonstrate the effectiveness of Bi-CoG, which consistently and significantly improves the performance of existing methods.
Introduction
@@
Addressing downstream tasks with limited labeled data remains a critical challenge in real-world applications (Yang et al. 2022). Two prominent paradigms have emerged to tackle this issue: (1) adapting pre-trained models, such as vision-language models (VLMs) (Radford et al. 2021; Jia et al. 2021), for image classification tasks; and (2) applying semi-supervised learning (SSL) methods that leverage large amounts of inexpensive unlabeled data. Recent work (Lv et al. 2025b) has further suggested combining these two approaches, where the strong zero-shot capabilities of pre-trained VLMs are used to generate reliable pseudo-labels that guide semi-supervised fine-tuning, thereby fully exploiting the strengths of both paradigms.
@@
Figure 1: Comparison of existing pseudo-labeling paradigms and our proposed Bi-CoG. (a) Majority voting from multiple VLMs; (b) Top-$k$ selection based on predicted logits; (c) Bi-CoG dynamically generate accurate and unbiased pseudo-labels based on bi-consistency and model error.
@@
Recent efforts have explored integrating these two paradigms by leveraging pre-trained models to extract domain knowledge from unlabeled data for downstream task adaptation. A common approach is model ensembling (Figure 1.a), such as V-PET (Gupte et al. 2024), which aggregates multi-view one-hot predictions from multiple vision models. As shown in Figure 1.b, another line of work selects top-$k$ predictions using class-wise or confidence-based thresholds. For instance, GRIP (Menghini, Delworth, and Bach 2023) and UPL (Huang, Chu, and Wei 2022) enforce class balance by selecting equal samples per class; CPL (Zhang et al. 2024) refines labels via instance-level confidence matrices to reduce pre-training bias.
@@
However, these methods struggle to balance the trade-off between pseudo-label accuracy and bias. Ensemble-based approaches can produce highly accurate pseudo-labels, but they often inadvertently amplify biases inherited from pre-training, leading to uneven performance across subgroups—a phenomenon commonly referred to as the “Matthew effect” (Zhu, Luo, and Liu 2022; Chen et al. 2022; Gan and Wei 2024). In contrast, soft-labeling techniques and class-wise pseudo-label allocation strategies can alleviate overconfidence to some extent, but they tend to introduce additional label noise and are highly sensitive to manually tuned hyperparameters. To the best of our knowledge, adaptively generating reliable and unbiased pseudo-labels in response to model performance remains a critical yet underexplored challenge in semi-supervised fine-tuning.
@@
As a plug-and-play approach, Bi-CoG can be seamlessly integrated into existing VLM fine-tuning pipelines. To evaluate its effectiveness, we conduct experiments under both open-world generalization (Zhou et al. 2024) and standard semi-supervised learning(Yang et al. 2022) settings, combining Bi-CoG with various prompt-tuning methods. Extensive results demonstrate the effectiveness of our bi-consistency-guided dynamic self-training strategy, which consistently improves performance across 14 datasets without relying on larger models or external knowledge, outperforming state-of-the-art (SOTA) methods.
@@
  • We design a deeply coupled self-training strategy integrates inter-model consistency, intra-model consistency and model error estimation, enabling reliable, low-bias pseudo-label generation and adaptive training.
@@
While pre-trained VLMs demonstrate strong zero-shot generalization in domain-specific downstream tasks, their large parameter sizes often hinder efficient adaptation to specific applications (Zhou et al. 2022b). To address this, recent works have proposed parameter-efficient fine-tuning approaches that introduce a small number of trainable parameters while keeping the majority of the model frozen (Li et al. 2024). Tip-Adapter (Zhang et al. 2022) and CLIP-Adapter (Gao et al. 2024) append lightweight bottleneck layers after the text or image encoders to learn residual features for better representation. CoOp (Zhou et al. 2022b) introduces prompt tuning to VLMs for the first time by replacing manually crafted hard prompts (e.g., “a photo of a classname”) with learnable soft text prompts optimized via backpropagation. Building on this idea, methods such as MaPLe (Khattak et al. 2023a), PromptSRC (Khattak et al. 2023b), and BMIP (Lv et al. 2025a) extend prompt tuning to both visual and textual encoders, enabling joint vision-language adaptation. While these methods have advanced the application of VLMs to downstream tasks, the transformation of prompts into continuous vectors makes them prone to overfitting and leads to sub-optimal performance in open-world scenarios (Zhou et al. 2022a; Khattak et al. 2023b).
@@
Semi-supervised learning (SSL) addresses label scarcity by leveraging both labeled and unlabeled data. Pseudo-labeling, the most widely used SSL approach, assigns labels to unlabeled data based on model predictions and improves performance through self-training. The key to effective pseudo-labeling lies in the generation of high-quality pseudo-labels (Yang et al. 2022). Traditional SSL methods typically filter pseudo-labels using confidence thresholds or adopt data augmentation to generate diverse views, aiming to improve label quality and model robustness (Lee et al. 2013a). For VLMs, their strong generalization ability allows them to produce highly accurate pseudo-labels even in zero- or few-shot settings, making them strong candidates as pseudo-labelers. Recent works have explored more advanced strategies: GRIP (Menghini, Delworth, and Bach 2023) gradually increases the upper limit on the number of pseudo-labels to mitigate calibration errors (Oh et al. 2024) and label imbalance (Wang et al. 2022); and CPL (Zhang et al. 2024) progressively refines pseudo-labels based on a confidence score matrix to avoid overconfidence. Despite their effectiveness, these methods struggle with balancing pseudo-label noise and inherited pre-training biases (Oh et al. 2024; Wang et al. 2022), while failing to dynamically adapt to model performance, ultimately limiting their potential.
@@
Figure 2: The overall framework of Bi-CoG’s pseudo-label selection. During the self-training phase, Bi-CoG employs inter-model consistency, intra-model consistency, and error-aware filtering to generate accurate and unbiased pseudo-labels, which are then used to fine-tune the target VLM.
Preliminary
In this section, we first revisit the core principles of CLIP (Radford et al. 2021), one of the most widely used VLMs, along with its associated prompt tuning methods, and subsequently formalize our problem definition.
CLIP.
CLIP comprises a text encoder $\psi$ and an image encoder $\phi$, aligned via pretraining on massive image-text pairs. Given an image $x$ and a set of class labels $\mathcal{Y}={y_{1},\dots,y_{C}}$, CLIP classifies $x$ by computing the similarity between the image representation $z=\phi(x)$ and each class representation $w_{i}=\psi(\pi_{i})$, where $\pi_{i}$ is a natural language prompt such as "a photo of a [CLASSi]". The predicted probability is computed as follows:
$p(y=i\mid x)=\frac{\exp(\cos(\phi(x),\psi(w_{i}))/\tau)}{\sum_{j=1}^{C}\exp(\cos(\phi(x),\psi(w_{j}))/\tau)}$ (1)
where $\tau$ is a temperature parameter that controls the sharpness of the probability distribution.
Prompt Tuning.
As a representative PEFT method for CLIP-based VLMs, prompt tuning learns adaptable prompts instead of fine-tuning the text encoder $\psi$. For instance, some studies (Zhou et al. 2022b, a) replace manually designed prompt templates with learnable prompt vectors $P=[t_{1},t_{2},\dots,t_{n}]$, where each $t_{i}$ is a trainable token vector and $n$ denotes the prompt length. Given a set of class names $c_{1},c_{2},\dots,c_{N}]$, the prediction probability is computed as:
$p(y=i\mid x)=\frac{\exp(\cos(\phi(x),\psi([P,c_{i}]))/\tau)}{\sum_{j=1}^{N}\exp(\cos(\phi(x),\psi([P,c_{j}]))/\tau)}$ (2)
where $\cdot$$\cdot$] denotes concatenation. The prompt $P$ is optimized via backpropagation during fine-tuning.
Problem Definition.
@@
In downstream scenarios with scarce labeled data, we typically have access to a small labeled dataset denoted as $D_{L}=\{(x_{i},y_{i})\}_{i=1}^{M}$, along with a large amount of unlabeled data $D_{UL}=\{x_{i}\}_{i=1}^{N}$, where $M\ll N$. Given a collection of $K$ pre-trained VLMs $\{f_{k}\}_{k=1}^{K}$ (with $K\geq 3$), the objective is to effectively leverage both $D_{L}$ and $D_{UL}$ to fine-tune the models and maximize performance on the target dataset $D_{test}$.
Methodology
@@
The main procedure of Bi-CoG is summarized in Algorithm 1, and we elaborate on each component in the subsequent sections.
Inter-model Consistency
During self-training with pseudo-labeled data, label fidelity is critical to effective and stable optimization (Angluin and Laird 1988). Excessive noise can mislead learning, causing convergence to suboptimal or incorrect representations. To mitigate this issue, we first generate pseudo-labels based on inter-model consistency by integrating predictions of multiple VLMs, with the primary goal of ensuring the accuracy of pseudo-labels.
@@
where $\mathcal{Y}$ denotes the label space, and $\mathbb{I}[\cdot]$ is the indicator function. This inter-model consistency strategy ensembles multiple models to generate more reliable pseudo-labels, reducing noise from individual prediction errors. Moreover, the leave-one-out scheme prevent amplifying biases inherent to the target model, thereby resulting in more accurate and unbiased candidate datasets for fine-tuning.
Intra-model Consistency
While inter-model consistency improves pseudo-label reliability by aggregating agreement across models, it may still preserve instances where VLMs share similar biases, often due to overlapping pre-training corpora or architectural homogeneity (Gan and Wei 2024; Wang and Russakovsky 2023). To further mitigate biased supervision, we introduce intra-model consistency, which leverages the consistency of a single model’s predictions under different input views to refine the pseudo-labeled sets obtained in the previous stage.
@@
Once the $D_{Intra}^{j}$ are computed, the candidate pseudo-label set for target model $f_{j}$ is constructed by intersecting its inter-model consistent set $D_{Inter}^{j}$ with the intra-model consistent set :
@@
The intersection of pseudo-labels from inter- and intra-model consistency preserves accuracy while mitigating shared model biases. Consequently, the resulting pseudo-labeled datasets are both highly reliable and low in bias, providing a robust foundation for the subsequent dynamic self-training stage.
Error-aware Filtering
After bi-consistency-guided pseudo-labeling, we obtain a candidate sample set $D_{PL}$, which contrains high-accuracy and low-biased pseudo-labels. However, during self-training, it is crucial to determine an appropriate upper bound on the number of pseudo-labeled samples based on model performance. Prior studies (Angluin and Laird 1988; Chen et al. 2022) have shown that even minor label noise, if not aligned with model competence, can significantly degrade performance, making further noise reduction in $D_{PL}$ essential. Existing methods (Menghini, Delworth, and Bach 2023; Huang, Chu, and Wei 2022) typically rely on fixed thresholds to limit the number of pseudo-labels, which may be suboptimal as model performance evolves over time.
To address this limitation, we dynamically adjust the upper bound on pseudo-label usage based on estimated model error ratio, guided by a theoretical analysis. Specifically, following prior work (Angluin and Laird 1988; Zhou and Li 2005) we formalize the influence of noisy supervision on model $f_{j}$ during iterative self-training:
@@
In practice, we estimate the theoretical error rate ratio from the performance of the K1K-1 models on the labeled set. However, this estimation can be biased, as repeated training on labeled data tends to produce stable—and often low—error ratios, which are reflected in Equation 7 as values close to 1. In contrast, the actual error ratio on unlabeled data typically decreases during training, making the true ratio $\frac{\hat{e}_{t}}{\hat{e}_{t-1}}$ lower than the estimated $\frac{e_{t}}{e_{t-1}}$. Ignoring this discrepancy would introduce bias into the estimated pseudo-label budget, compromising the validity of the theoretical analysis. To address this issue, we introduce Theorem 1 to provide a correction:
@@
A detailed proof of Lemma 1 and Theorem 1 is provided in the supplementary material. Based on the theoretical analysis, we establish the necessary condition under which the model can be effectively optimized by self-training, as described in Equation 9, and establish an upper bound on the pseudo-label budget for each iteration. Accordingly, the final set of pseudo-labeled samples is formalized as:
@@
In summary, Bi-CoG’ offers the following advantages:
@@
    Plug-and-Play Adaptability: Bi-CoG can be seamlessly integrated into existing pretrain-finetune pipelines, providing consistent performance improvements with minimal changes to the original framework.
@@
To answer RQ1 and RQ2, we evaluate the effectiveness of Bi-CoG on two widely adopted settings: open-world generation setting (Zhou et al. 2024; Lv et al. 2025a) and semi-supervised setting (Lee et al. 2013b; Yang et al. 2022). In open-world generation setting, the model is trained only on a subset of categories (referred to as base classes), while the remaining categories are treated as novel classes. We report the harmonic mean (HM) of the model’s performance on base and novel classes, as well as the overall accuracy on the full test set, which includes both base and novel classes. Semi-supervised setting follows a common protocol where only a small portion of the data is labeled, while the majority remains unlabeled.
@@
We evaluate the performance of our method on 14 recognition datasets, including 11 commonly used for assessing VLMs and 3 classical benchmarks for semi-supervised learning. Detailed information about the datasets is provided in the supplementary material. We select CLIP (Radford et al. 2021) and its three representative prompt tuning methods as baselines: CoOp (Zhou et al. 2022b), MaPLe (Khattak et al. 2023a), and PromptSRC (Khattak et al. 2023b), among which PromptSRC represents the current SOTA method. For methods that integrate the two paradigms, we select GRIP (Menghini, Delworth, and Bach 2023) and CPL (Zhang et al. 2024) as baselines, with CPL representing the current SOTA approach. The parameters of Bi-CoG and additional experimental details are provided in the supplementary material.
Experimental Results

### v3

@@ Abstract
Exploiting unlabeled data through semi-supervised learning (SSL) or leveraging pre-trained models via fine-tuning are two prevailing paradigms for addressing label-scarce scenarios. Recently, growing attention has been given to combining fine-tuning of pre-trained vision-language models (VLMs) with SSL, forming the emerging paradigm of semi-supervised fine-tuning. However, existing methods often suffer from model bias and hyperparameter sensitivity, due to reliance on prediction consistency or pre-defined thresholds. To address these limitations, we propose a simple yet effective plug-and-play method named Bi-Consistency-Guided Self-Training (Bi-CoG), which assigns accurate and low-bias pseudo-labels, by simultaneously exploiting inter- and intra-model consistency, along with an error-aware dynamic filtering strategy. Both theoretical analysis and extensive experiments over 14 datasets demonstrate the effectiveness of Bi-CoG, which consistently and significantly improves the performance of existing methods. Codes are available at   Bi-CoG.
11footnotetext: Equal contribution.22footnotetext: Corresponding author.
1 Introduction
@@
Addressing downstream tasks with limited labeled data remains a critical challenge in real-world applications Yang et al. (2023); Shao et al. (2026); Yang et al. (2025). Two prominent paradigms have emerged to tackle this issue: (1) adapting pre-trained models, such as vision-language models (VLMs) Radford et al. (2021); Jia et al. (2021), for downstream tasks; and (2) applying semi-supervised learning (SSL) methods that leverage large amounts of inexpensive unlabeled data. Recent work Lv et al. (2025b) has further suggested combining them, where the strong zero-shot capabilities of pre-trained VLMs are used to generate reliable pseudo-labels that guide semi-supervised fine-tuning, thereby fully exploiting the strengths of both paradigms.
@@
Figure 1: Comparison of existing pseudo-labeling paradigms and Bi-CoG. (a) Majority voting from multiple VLMs; (b) Top-$k$ selection based on predicted logits; (c) Bi-CoG dynamically generates pseudo-labels based on bi-consistency and model error.
@@
Recent efforts have explored integrating these two paradigms by leveraging pre-trained models to extract domain knowledge from unlabeled data for downstream task adaptation. A common approach is model ensembling (Figure 1.a), such as V-PET Gupte et al. (2024), which aggregates multi-view one-hot predictions from multiple vision models. As shown in Figure 1.b, another line of work selects top-$k$ predictions using class-wise or confidence-based thresholds. For instance, GRIP Menghini et al. (2023) and UPL Huang et al. (2022) enforce class balance by selecting equal samples per class; CPL Zhang et al. (2024) refines labels via instance-level confidence matrices.
@@
However, these methods struggle to balance the trade-off between pseudo-label accuracy and bias. Ensemble-based approaches can produce highly accurate pseudo-labels, but they often inadvertently amplify biases inherited from pre-training, leading to uneven performance across subgroups—a phenomenon commonly referred to as the “Matthew effect” Zhu et al. (2022); Chen et al. (2022); Gan and Wei (2024); Guo et al. (2020). In contrast, soft-labeling techniques and class-wise pseudo-label allocation strategies can alleviate overconfidence to some extent, but they tend to introduce additional label noise and are highly sensitive to manually tuned hyperparameters. To the best of our knowledge, adaptively generating reliable and unbiased pseudo-labels in response to model performance remains a critical yet underexplored challenge in semi-supervised fine-tuning.
@@
As a plug-and-play approach, Bi-CoG can be seamlessly integrated into existing VLM fine-tuning pipelines. To evaluate its effectiveness, we conduct experiments under both open-world generalization Zhou et al. (2024) and standard semi-supervised learningYang et al. (2023) settings, combining Bi-CoG with various prompt-tuning methods. Extensive results demonstrate the effectiveness of our bi-consistency-guided dynamic self-training strategy, which consistently improves performance across 14 datasets without relying on larger models or external knowledge, outperforming state-of-the-art (SOTA) methods.
@@
  • We design a deeply coupled self-training strategy that integrates inter-model consistency, intra-model consistency and model error estimation, enabling reliable, low-bias pseudo-label generation and adaptive training.
@@ 2 Related Work
While pre-trained VLMs demonstrate strong zero-shot generalization in domain-specific downstream tasks, their large parameter sizes often hinder efficient adaptation to specific applications Zhou et al. (2022b). To address this, recent works have proposed parameter-efficient fine-tuning approaches that introduce a small number of trainable parameters while keeping the majority of the model frozen Li et al. (2024). Tip-Adapter Zhang et al. (2022) and CLIP-Adapter Gao et al. (2024) append lightweight bottleneck layers after the text or image encoders to learn residual features for better representation. CoOp Zhou et al. (2022b) introduces prompt tuning to VLMs for the first time by replacing manually crafted hard prompts (e.g., “a photo of a classname”) with learnable soft text prompts optimized via backpropagation. Building on this idea, methods such as MaPLe Khattak et al. (2023a), PromptSRC Khattak et al. (2023b), and BMIP Lv et al. (2025a) extend prompt tuning to both visual and textual encoders, enabling joint vision-language adaptation. While these methods have advanced the application of VLMs to downstream tasks, the transformation of prompts into continuous vectors makes them prone to overfitting and leads to sub-optimal performance in open-world scenarios Zhou et al. (2022a); Khattak et al. (2023b). Beyond prompt tuning, recent work on VLM selection and reuse Tan et al. (2025) further explores how to choose and adapt pre-trained models for downstream tasks.
@@
Semi-supervised learning (SSL) addresses label scarcity by leveraging both labeled and unlabeled data. Pseudo-labeling, the most widely used SSL approach, assigns labels to unlabeled data based on model predictions and improves performance through self-training. The key to effective pseudo-labeling lies in the generation of high-quality pseudo-labels Yang et al. (2023). Traditional SSL methods typically filter pseudo-labels using confidence thresholds or adopt data augmentation to generate diverse views, aiming to improve label quality and model robustness Lee and others (2013). For VLMs, their strong generalization ability allows them to produce highly accurate pseudo-labels even in zero- or few-shot settings, making them strong candidates as pseudo-labelers. Recent works have explored more advanced strategies: GRIP Menghini et al. (2023) gradually increases the upper limit on the number of pseudo-labels to mitigate calibration errors Oh et al. (2024) and label imbalance Wang et al. (2022); Guo and Li (2022); and CPL Zhang et al. (2024) progressively refines pseudo-labels based on a confidence score matrix to avoid overconfidence. Despite their effectiveness, these methods struggle with balancing pseudo-label noise and inherited pre-training biases Oh et al. (2024); Wang et al. (2022); Li et al. (2021b), while failing to dynamically adapt to model performance in open environments Guo et al. (2025), ultimately limiting their potential.
@@
Figure 2: The overall framework of Bi-CoG’s pseudo-label selection. During the self-training phase, Bi-CoG employs inter-model consistency, intra-model consistency, and error-aware filtering to generate accurate and unbiased pseudo-labels.
Problem Definition
@@
Based on the aforementioned challenges in pseudo-labeling—specifically the issues of label noise and static adaptation—we formally define our task. In downstream scenarios with scarce labeled data, we typically have access to a small labeled dataset denoted as $D_{L}=\{(x_{i},y_{i})\}_{i=1}^{M}$, along with a large amount of unlabeled data $D_{UL}=\{x_{i}\}_{i=1}^{N}$, where $M\ll N$. Given a collection of $K$ pre-trained VLMs $\{f_{k}\}_{k=1}^{K}$ (with $K\geq 3$), the objective is to effectively leverage both $D_{L}$ and $D_{UL}$ to fine-tune the models and maximize performance on the target dataset $D_{test}$.
3 Methodology
@@
The procedure of Bi-CoG is summarized in Algorithm 1. We elaborate on each component in the subsequent sections.
3.1 Inter-model Consistency
During self-training with pseudo-labels, label fidelity is critical to effective and stable optimization Angluin and Laird (1988). Excessive noise can mislead learning, causing convergence to suboptimal representations. To mitigate this issue, we first generate pseudo-labels based on inter-model consistency by integrating predictions of multiple VLMs, with the primary goal of ensuring the accuracy of pseudo-labels.
@@ 3.1 Inter-model Consistency
where $\mathcal{Y}$ denotes the label space, and $\mathbb{I}[\cdot]$ is the indicator function. This strategy ensembles multiple models to generate reliable pseudo-labels, reducing noise from individual prediction errors. Moreover, the leave-one-out scheme prevents amplifying biases inherent to the target model, thereby resulting in more accurate and unbiased data for fine-tuning.
3.2 Intra-model Consistency
While inter-model consistency improves pseudo-label reliability by aggregating agreement across models, it may still preserve instances where VLMs share similar biases, often due to overlapping pre-training corpora or architectural homogeneity Gan and Wei (2024); Wang and Russakovsky (2023). To further mitigate biased supervision, we introduce intra-model consistency, which leverages the consistency of a single model’s predictions under different input views to refine the pseudo-labeled sets obtained in the previous stage.
@@ 3.2 Intra-model Consistency
Once the $D_{Intra}^{j}$ are computed, the candidate pseudo-label set is constructed by intersecting its inter-model consistent set $D_{Inter}^{j}$ with the intra-model consistent set:
@@
The intersection of pseudo-labels from inter-model consistency and intra-model consistency preserves accuracy while mitigating shared model biases. Consequently, the resulting pseudo-labeled datasets are both highly reliable and low in bias, providing a robust foundation for the subsequent dynamic self-training stage.
3.3 Error-aware Filtering
After bi-consistency-guided pseudo-labeling, we obtain a candidate sample set $D_{PL}$, which contains high-accuracy and low-biased pseudo-labels. However, during self-training, it is crucial to determine an appropriate upper bound on the number of pseudo-labeled samples based on model performance. Prior studies Angluin and Laird (1988); Chen et al. (2022) have shown that even minor label noise, if not aligned with model competence, can significantly degrade performance, making further noise reduction in $D_{PL}$ essential. Existing methods Menghini et al. (2023); Huang et al. (2022) typically rely on fixed thresholds to limit the number of pseudo-labels, which can be suboptimal as model performance evolves over time and across different datasets.
To address this limitation, we dynamically adjust the upper bound on pseudo-label usage based on estimated model error ratio, guided by a theoretical analysis. Specifically, following prior work Angluin and Laird (1988); Zhou and Li (2005) we formalize the influence of noisy supervision on model $f_{j}$ during iterative self-training:
@@ 3.3 Error-aware Filtering
In practice, we estimate the theoretical error rate ratio from the performance of the K1K-1 models on the labeled set. However, this estimation can be biased, as repeated training on labeled data tends to produce stable—and often low—error ratios, which are reflected in Equation 5 as values close to 1. In contrast, the actual error ratio on unlabeled data typically decreases during training, making the true ratio $\frac{e_{t}}{e_{t-1}}$ lower than the estimated $\frac{\hat{e}_{t}}{\hat{e}_{t-1}}$. Ignoring this discrepancy would introduce bias into the estimated pseudo-label budget, compromising the validity of the theoretical analysis. To address this issue, we introduce Theorem 1 to provide a correction:
@@
Based on the theoretical analysis, we establish the necessary condition under which the model can be effectively optimized by self-training, as described in Equation 7, and establish an upper bound on the pseudo-label budget for each iteration. Here, $L_{j}^{\prime}$ denotes the size of the pseudo-labeled training set used by model $f_{j}$ in the previous iteration. Accordingly, the final set of pseudo-labeled samples is formalized as:
@@
After training, Bi-CoG adopts a majority voting strategy to generate predicted labels for evaluation during the testing phase, ensuring robust and reliable assessment results.
In summary, Bi-CoG offers the following key advantages:
@@
    Plug-and-Play Adaptability: Bi-CoG can be seamlessly integrated with existing pre-trained models, providing consistent performance improvements with minimal changes to the original method.
@@ 4.1 Experimental Settings
To answer RQ1 and RQ2, we evaluate the effectiveness of Bi-CoG on two widely adopted settings: open-world generalization setting Zhou et al. (2024); Lv et al. (2025a) and semi-supervised setting Lee and others (2013); Yang et al. (2023). In open-world generalization setting, the model is trained only on a subset of categories (referred to as base classes), while the remaining categories are treated as novel classes. We report the harmonic mean (HM) of the model’s performance on base and novel classes, as well as the overall accuracy on the full test set, which includes both base and novel classes. Semi-supervised setting follows a common protocol where only a small portion of the data is labeled, while the majority remains unlabeled.
@@
We evaluate the performance of our method on 14 recognition datasets, including 11 commonly used for assessing VLMs and 3 classical benchmarks for semi-supervised learning. We select CLIP Radford et al. (2021) and its three representative prompt tuning methods as baselines: CoOp Zhou et al. (2022b), MaPLe Khattak et al. (2023a), and PromptSRC Khattak et al. (2023b), among which PromptSRC represents the current SOTA method. For methods that integrate the two paradigms, we select GRIP Menghini et al. (2023) and CPL Zhang et al. (2024) as baselines, with CPL representing the current SOTA approach.
Implementation Details.
We set the number of VLMs $K=3$ throughout all experiments. The setting $K\geq 3$ ensures that an absolute majority can be achieved in most cases, facilitating reliable pseudo-label selection. While a larger $K$ further improves stability, it also increases GPU memory consumption. We adopt $K=3$ to balance effectiveness and efficiency. For the scaling factor $\alpha$, it provides a controllable mechanism to adjust the strictness of pseudo-label filtering: a smaller $\alpha$ enhances the utilization of more pseudo-labels, while a larger $\alpha$ suppresses noisy pseudo-labels by raising the filtering threshold. We uniformly set $\alpha=1$ across all datasets to ensure generalizability without dataset-specific tuning. All experiments are conducted on a single NVIDIA A800 GPU.
4.2 Experimental Results